HELP!!! PLEASE!!!!
(3x^2 under the radical sign) + (2 outside radical sign 48x^2 under the radical sign) + (3 outside radical sign 48x^2 under the radical sign)

Answer by uma(370)

The question is :
sqrt(3x^2) + 2sqrt(48x^2) + 3sqrt(48x^2)
= sqrt(3x^2) + 5sqrt(48x^2) [adding the last two terms]
= sqrt(3)*sqrt(x^2) + 5sqrt(48) * sqrt(x^2)
= sqrt(3) * x + 5sqrt(16*3) * x
= xsqrt(3) + 5x sqrt(16)*sqrt(3)
= xsqrt(3) + 5x* 4 * sqrt(3)
= x sqrt(3) + 20x sqrt(3)
= 21x sqrt(3)


Good Luck!!!